# DDAT: Diffusion Policies Enforcing Dynamically Admissible Robot Trajectories

### v1

@@ VII-C Training with projections
Let us now answer Q3 on whether incorporating projections during training help diffusion models generate better samples. Our intuition was that projections during training would better prepare the diffusion models for the projections at inference. However, Figures 5(b) and 5(c) both show little difference in trajectory quality between the models trained with projections and those using projections only at inference. The only significant quality difference is caused by employing different projectors. In terms of admissibility, Figures 7(b), 7(c), 7(e), and 7(f) show no significant difference between models trained with projections and those only using them at inference. Despite our best effort to incorporate projections into the training of diffusion models, projections still disrupt training by significantly increasing the training loss. Thus our answer to Q3 is negative, training with projections is not helpful.
@@
TABLE III: Mean and standard deviation of reward and survival rates (%) over 100 initial states with 8 trajectories sampled for each. Among all the models tested $SA\mathcal{P}^{\text{SA}}_{\sigma}$ consistently perform among the best.
@@ VII-D Diffusion modality
D-MPC [52] is an offline model-based approach learning a dynamics model and a policy using two different diffusion models. Because our work assumes access to a dynamics simulator, a fair comparison with D-MPC should only compare its diffusion policy, a conditional model predicting actions conditioned on the current state. We trained such a model on our datasets using our DiT architecture under the acronym $A(\cdot|s_{0})$.
@@
As shown on Table V the Action model performance can vary from worse to best on extremely similar environments. Except for the Hopper where $A(\cdot|s_{0})$ performs poorly, it shows decent results on the other environments of Table IV. Due to its inconsistency predicting only action sequences does not perform as well as state-action models like $SA\mathcal{P}_{\sigma}^{\text{SA}}$, which answers Q4.
@@
TABLE V: Reward and survival rates (%) ±$\pm std$ over 100 initial states $s_{0}$ with 64 samples generated per $s0$. All these models predict actions implemented open-loop on the robots and conditioned on the velocity command $c$. We could not evaluate state models for the Unitree due to the lack of performant inverse dynamics model. The action only mnodel $A(\cdot|c,s_{0})$ ranks respectively best and worse on the GO1 and GO2, whereas state-action model $SA\mathcal{P}^{\text{A}}_{\sigma}(\cdot|c)$ performs consistently well on both robots.
@@ VIII Hardware Experiments
In this section, we evaluate DDAT on real-world hardware. Specifically, we demonstrate that DDAT can generate high-dimensional trajectories deployable on hardware in open-loop. Moreover, DDAT causes fewer failures than an baseline diffusion model and outperforms it in task completion. We train both diffusion policies in the Unitree GO2 environment and generate trajectories walking forward for $10$ seconds. To evaluate walking behavior, we aim for the quadruped to pass a goal line $3$ meters in front of the start position while staying within a $1$ meter corridor. If the quadruped leaves the sidelines or begins to fall, we consider it a failure. For this evaluation, we sampled $3$ trajectories from each model and performed $10$ trials. As such, we performed a total of $30$ trials for each model, and a summary of the results is presented in Table VI.
@@
Overall, we found that DDAT had consistently higher success rates compared to the vanilla diffusion policy. Specifically, DDAT was successful more than twice as many trials as the diffusion policy baseline. The baseline would often generate actions causing the robot to slip or begin to fall, leading to either a ‘Fall Failure’ or a recovery performed by the onboard GO2 computer. However, even after such a recovery, the quadruped would often be pointed in the wrong direction, leading to the robot to leave the our prescribed corridor, making it a ‘Exit Region Failure’. The baseline trajectory would also occasionally cause the quadruped motors to vibrate and overwork themselves until they locked up, and the robot would begin to slouch in place. We also considered such scenarios as ‘Fall Failure’. In comparison, our DDAT trajectories were often smoother, commanding the Unitree GO2 to follow a periodic motion, leading it along a fast and convenient path to the goal line at $3$ meters without much deviation. Nonetheless, as these are open loop trajectories, the stochastic nature of the physical hardware caused some of the trajectories to leave the region. Overall, DDAT’s superior performance showcases the importance of developing dynamically feasible and admissible trajectories for real world deployments.
@@ IX Limitations
Another limitation shared by most diffusion policies is their slow and computationally intensive inference. Compared to the traditional DDPM [18] diffusion models requiring hundreds of inference steps, our implementation only requires a fraction of the inference steps (we used $N=5$) but the projections can become costly for long-horizon and large action spaces. Speed improvements can certainly be achieved following an approach similar to that of [19].
@@ Appendix C Diffusion implementation details
TABLE VII: Environments and DiT parameters. Each DiT model generates either sequences of states ”S”, states and actions ”SA”, or only actions ”A”. The Unitree models are all conditioned on the velocity command $c$.
@@ D-B Quadcopter
Its state is composed of the position of its center of mass $x,y,z$, quaternion orientation $q_{0},q_{1},q_{2},q_{3}$, linear velocity of the center of mass $v_{x},v_{y},v_{z}$, angular velocity $p,q,r$ and angular velocities of the propellers $w_{1},w_{2},w_{3},w_{4}$. The actions are the angular acceleration of the propellers $\dot{w}_{1},\dot{w}_{2},\dot{w}_{3},\dot{w}_{4}$. The quadcopter starts around the origin at hover and is tasked with reaching position $x=7$ in minimum time thanks to a model predictive controller (MPC) implemented in CasADi [3]. The quadcopter must dodge two circular obstacles shown on Fig. 8(c). To reach the target in minimal time while dodging the obstacles, the MPC takes full advantage of all the degrees of freedom of the quadcopter as seen on Fig. 8(a) and pushes the propellers to their maximal speed as shown on Fig. 8(b).
@@
Figure 8: Quadcopter reference trajectory generated using MPC [44].
@@ D-C Unitree GO1
For the training dataset, we trained a joystick policy that can follow random commands consisting of the linear velocity of the base $v_{x}$, $v_{y}$, and yaw velocity of base $r$ using MuJoCo playground [50]. We generated 1000 trajectories with randomized commands and train a diffusion model conditioned on commands. Each trajectory includes 500 steps. State space for DDAT consists of the position of the base $x,y,z$, quaternion orientation of the base $q_{0},q_{1},q_{2},q_{3}$, linear velocity of the base $v_{x},v_{y},v_{z}$, linear acceleration of the base $a_{x},a_{y},a_{z}$, angle of each joint $\theta_{1},\dots,\theta_{12}$, and angular velocity of each joint $\dot{\theta}_{1},\dots,\dot{\theta}_{12}$. Action space is the target joint angles $\theta_{1}^{*},\dots,\theta_{12}^{*}$ for the low-level PD controller.
No inverse dynamics as Algorithms 6 and 7 could not achieve sufficient precision in a reasonable time, and its dynamics is a black-box since simulated with MuJoCo [50].
@@ D-D Unitree GO2
As Unitree GO2 simulation is not supported by MuJoCo playground [50], we implement a custom environment based on GO1 environment and MuJoCo menagerie [49]. We use the same state space and action space as the GO1 environment for the diffusion planner and train it with 1000 trajectories with randomized commands.
Similar to the GO1 environment, no inverse dynamics as Algorithms 6 and 7 could not achieve sufficient precision in a reasonable time, and its dynamics is a black-box since simulated with MuJoCo [43].

### v2

@@ VII-C Training with projections
Let us now answer Q3 on whether incorporating projections during training helps diffusion models generate better samples. Our intuition was that projections during training would better prepare the diffusion models for the projections at inference. However, Figures 5(b) and 5(c) both show little difference in trajectory quality between the models trained with projections and those using projections only at inference. The only significant quality difference is caused by employing different projectors. In terms of admissibility, Figures 7(b), 7(c), 7(e), and 7(f) show no significant difference between models trained with projections and those only using them at inference. Despite our best effort to incorporate projections into the training of diffusion models, projections still disrupt training by significantly increasing the training loss. Thus our answer to Q3 is negative, training with projections is not helpful.
@@
TABLE III: Mean and standard deviation of reward and survival rates (%) over 100 initial states with 8 trajectories sampled for each. As in [20] we used Model Predictive Path Integral (MPPI) [48] as a baseline, but it struggled due to the long-horizon nature of our tasks. The data row refers to our training data generated by closed-loop RL or MPC policies, which is why their performance exceeds all the open-loop models. Among all the models tested $SA\mathcal{P}^{\text{SA}}_{\sigma}$ consistently perform among the best.
@@ VII-D Diffusion modality
D-MPC [54] is an offline model-based approach learning a dynamics model and a policy using two different diffusion models. Because our work assumes access to a dynamics simulator, a fair comparison with D-MPC should only compare its diffusion policy. We trained such a model generating sequences of actions conditioned on the initial state $s_{0}$ on our datasets using our DiT architecture under the acronym $A(\cdot|s_{0})$.
@@
As shown in Table V the Action model performance can vary from worse to best on extremely similar environments. Except for the Hopper where $A(\cdot|s_{0})$ performs poorly, it shows decent results on the other environments of Table IV. Due to its inconsistency predicting only action sequences does not perform as well as state-action models like $SA\mathcal{P}_{\sigma}^{\text{SA}}$, which answers Q4.
@@
TABLE V: Reward and survival rates (%) ±$\pm std$ over 100 initial states $s_{0}$ with 64 samples generated per $s_{0}$. All these models predict actions implemented open-loop on the robots and conditioned on the velocity command $c$. We could not evaluate state models for the Unitree due to the lack of performant inverse dynamics model. The action only model $A(\cdot|c,s_{0})$ ranks respectively best and worse on the GO1 and GO2, whereas state-action model $SA\mathcal{P}^{\text{A}}_{\sigma}(\cdot|c)$ performs consistently well on both robots.
@@ VIII Hardware Experiments
In this section, we evaluate DDAT on real-world hardware. Specifically, we demonstrate that DDAT can generate high-dimensional trajectories deployable on hardware in open-loop. Moreover, DDAT causes fewer failures than a baseline diffusion model and outperforms it in task completion. We train both diffusion policies in the Unitree GO2 environment and generate trajectories walking forward for $10$ seconds. To evaluate walking behavior, we aim for the quadruped to pass a goal line $3$ meters in front of the start position while staying within a $1$ meter corridor. If the quadruped leaves the sidelines or begins to fall, we consider it a failure. For this evaluation, we sampled $3$ trajectories from each model and performed $10$ trials. As such, we performed a total of $30$ trials for each model, and a summary of the results is presented in Table VI.
@@
Overall, we found that DDAT had consistently higher success rates compared to the vanilla diffusion policy. Specifically, DDAT was successful in more than twice as many trials as the diffusion policy baseline. The baseline would often generate actions causing the robot to slip or begin to fall, leading to either a ‘Fall Failure’ or a recovery performed by the onboard GO2 computer. However, even after such a recovery, the quadruped would often be pointed in the wrong direction, leading to the robot to leave the prescribed corridor, making it a ‘Exit Region Failure’. The baseline trajectory would also occasionally cause the quadruped motors to vibrate and overwork themselves until they locked up, and the robot would begin to slouch in place. We also considered such scenarios as ‘Fall Failure’. In comparison, our DDAT trajectories were often smoother, commanding the Unitree GO2 to follow a periodic motion, leading it along a fast and straight path to the goal line at $3$ meters without much deviation. Nonetheless, as these are open loop trajectories, the stochastic nature of the physical hardware caused some of the trajectories to leave the region. Overall, DDAT’s superior performance showcases the importance of developing dynamically feasible and admissible trajectories for real world deployments.
@@ IX Limitations
Another limitation shared by most diffusion policies is their slow and computationally intensive inference. Compared to the traditional DDPM [18] models requiring hundreds of inference steps, our implementation only requires a fraction of the inference steps (we used $N=5$) but the projections can become costly for long-horizon and large action spaces. Speed improvements can certainly be achieved following an approach similar to that of [19].
@@ X Conclusion and Future Work
Acknowledgments
This work is supported in part by the National Science Foundation, under grants ECCS-2438314 CAREER Award, CNS-2423130, CCF-2423131, and by the Google-BAIR Commons Project Year 6. This work is supported in part by the Robotics and AI Institute. K. Sreenath has financial interest in the Robotics and AI Institute. He and the company may benefit from the commercialization of the results of this research.
The authors are also grateful to John Viljoen for sharing the code of his MPC and quadcopter environment.
@@ Appendix C Diffusion implementation details
TABLE VII: Environments and DiT parameters. Each DiT model generates either sequences of states ”S”, states and actions ”SA”, or only actions ”A”. The Unitree models are all conditioned on the velocity command $cmd$.
@@ D-B Quadcopter
Its state is composed of the position of its center of mass $x,y,z$, quaternion orientation $q_{0},q_{1},q_{2},q_{3}$, linear velocity of the center of mass $v_{x},v_{y},v_{z}$, angular velocity $p,q,r$ and angular velocities of the propellers $w_{1},w_{2},w_{3},w_{4}$. The actions are the angular acceleration of the propellers $\dot{w}_{1},\dot{w}_{2},\dot{w}_{3},\dot{w}_{4}$. The quadcopter starts around the origin at hover and is tasked with reaching position $x=7$ in minimum time thanks to a model predictive controller (MPC) implemented in CasADi [3]. The quadcopter must dodge two circular obstacles shown in Fig. 8(c). To reach the target in minimal time while dodging the obstacles, the MPC takes full advantage of all the degrees of freedom of the quadcopter as seen in Fig. 8(a) and pushes the propellers to their maximal speed as shown in Fig. 8(b).
@@
Figure 8: Quadcopter reference trajectory generated using model predictive control (MPC) [45].
@@ D-C Unitree GO1
For the training dataset, we trained a joystick policy that can follow random commands consisting of the linear velocity of the base $v_{x}$, $v_{y}$, and yaw velocity of base $r$ using MuJoCo playground [52]. We generated 1000 trajectories with randomized commands and train a diffusion model conditioned on commands. Each trajectory includes 500 steps. The state space consists of the position of the base $x,y,z$, quaternion orientation of the base $q_{0},q_{1},q_{2},q_{3}$, linear velocity of the base $v_{x},v_{y},v_{z}$, linear acceleration of the base $a_{x},a_{y},a_{z}$, angle of each joint $\theta_{1},\dots,\theta_{12}$, and angular velocity of each joint $\dot{\theta}_{1},\dots,\dot{\theta}_{12}$. The action space is the target joint angles $\theta_{1}^{*},\dots,\theta_{12}^{*}$ for the low-level PD controller. We did not use any inverse dynamics as Algorithms 6 and 7 could not achieve sufficient precision in a reasonable time, and the GO1 dynamics are a black-box since simulated with MuJoCo [52].
@@ D-D Unitree GO2
As the Unitree GO2 simulation is not supported by MuJoCo playground [52], we implemented a custom environment based on GO1 environment and MuJoCo menagerie [51]. We used the same state space and action space as the GO1 environment for the diffusion planner and trained it with 1000 trajectories with randomized commands. Similar to the GO1 environment, we did not have access to any inverse dynamics as Algorithms 6 and 7 could not achieve sufficient precision in a reasonable time, and the GO2 dynamics are a black-box simulated with MuJoCo [44].
@@ D-E Hardware settings
D-F Computational overhead
The iterative dynamics sampling and the CVXPY optimization [1] of projectors (5) and (6) generate additional computational costs. Indeed, these approaches scales exponentially in the dimension of the action space. The action set $\mathcal{A}$ is typically a hypercube $[-1,1]^{m}$ with $2^{m}$ vertices. As shown on Table VIII, the models with reachable set projections $S\mathcal{P}_{\sigma}$ and $S\mathcal{P}_{\sigma}^{\text{ref}}$ are much slower than the model without projection $S$. That is why we introduced models $SA\mathcal{P}_{\sigma}^{\text{A}}$ and $SA\mathcal{P}_{\sigma}^{\text{SA}}$ capable of scaling to much larger action sets (like the GO2).
[…]
Hopper 0.1 28 7 0.2 0.5 1
Walker 0.3 32 20 0.5 0.8 1.3
Quadcopter 0.3 24 39 0.3 0.5 0.8
GO2 - - - 2.6 5.5 8.7
TABLE VIII: Inference time (in seconds) to generate 8 trajectories of 300 timesteps each on a cpu Intel Core i5-14400F 2.5GHz with 16GB of RAM. Our models $S\mathcal{P}_{\sigma}$ and $S\mathcal{P}_{\sigma}^{\text{ref}}$ are significantly slower than baseline $S$ due to their use of CVXPY projections and cannot scale up to the GO2. Our models $SA\mathcal{P}_{\sigma}^{\text{A}}$ and $SA\mathcal{P}_{\sigma}^{\text{SA}}$ are slower than the baseline $SA$ but scale better than the state models with CVXPY projections.
We believe our models with projections can be optimized to run faster following the method of [19] as for now no efforts were made to accelerate the computations. All the diffusion models were trained on a single Nvidia A100 gpu. Training took between a few hours for the Hopper to a full day for the Unitree GO2. We are now working on accelerating inference for closed-loop real time deployement.